# TSPRank: Bridging Pairwise and Listwise Methods with a Bilinear Travelling Salesman Model

### v2

@@ Abstract.
Traditional Learning-To-Rank (LETOR) approaches, including pairwise methods like RankNet and LambdaMART, often fall short by solely focusing on pairwise comparisons, leading to sub-optimal global rankings. Conversely, deep learning based listwise methods, while aiming to optimise entire lists, require complex tuning and yield only marginal improvements over robust pairwise models. To overcome these limitations, we introduce Travelling Salesman Problem Rank (TSPRank), a hybrid pairwise-listwise ranking method. TSPRank reframes the ranking problem as a Travelling Salesman Problem (TSP), a well-known combinatorial optimisation challenge that has been extensively studied for its numerous solution algorithms and applications. This approach enables the modelling of pairwise relationships and leverages combinatorial optimisation to determine the listwise ranking. TSPRank can be directly integrated as an additional component into embeddings generated by existing backbone models to enhance ranking performance. Our extensive experiments across three backbone models on diverse tasks, including stock ranking, information retrieval, and historical events ordering, demonstrate that TSPRank significantly outperforms both pure pairwise and listwise methods. Our qualitative analysis reveals that TSPRank’s main advantage over existing methods is its ability to harness global information better while ranking. TSPRank’s robustness and superior performance across different domains highlight its potential as a versatile and effective LETOR solution.
@@
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.14542444.
@@ 1. Introduction
Learning to Rank (LETOR) algorithms have become essential in applications such as recommendation systems (ada-ranker; peiPersonalizedRerankingRecommendation2019; rank-insights), question answering (10.1145/2505515.2505670; campese-etal-2024-pre), and information retrieval (aiLearningDeepListwise2018a; pang2020setrank; rankformer). These algorithms aim to order a list of ranking entities based on their features, optimising for the most relevant or preferred entities to appear at the top. Recently, LETOR methods have also expanded into other domains, such as stock and portfolio selection (SONG201720; portfolioselectionletor) and textual ordering (tan-etal-2013-learning; NEURIPS2023_f2a11632). Despite these broader applications of LETOR, latest fundamental work on LETOR still primarily focused on incorporating click data into ranking models and addressing biases introduced by user feedback (10.1145/3569453; 10.1145/3580305.3599914; 10.1145/3394486.3403336), concentrating on retrieval and recommendation tasks. Limited research has explored general ranking methods across diverse tasks and domains.
@@
Over the past decade, research on LETOR models mainly focuses on pairwise and listwise approaches, while pointwise methods often fail to capture the intricate inter-entity relationships that are essential for accurate ranking. Pairwise methods, such as LambdaMART and RankNet (lambdamart), primarily optimise for pairwise comparisons without a holistic view of the entire ranking list, potentially leading to sub-optimal global rankings. Listwise methods optimise the ranking of entire lists directly rather than individuals or pairs. This category includes advanced neural network architectures, particularly adaptations of the Transformer architecture (attention), such as Rankformer (rankformer) and SetRank (pang2020setrank). However, existing work shows that deep learning-based listwise models require complex tuning to achieve marginal gains over robust pairwise models like LambdaMART on information retrieval benchmarks (google-neural-outperform-gbdt). Therefore, pairwise and listwise methods each have inherent drawbacks and more effective solutions for LETOR have not been exhaustively explored.
@@
Recently, GNNRank successfully used directed graph neural networks to learn listwise rankings from pairwise comparisons (he2022gnnrank). Although their method is not applicable to general LETOR problems, as it requires pre-known pairwise relationships like the outcomes of sports matches, it still suggests a potential approach for LETOR by representing pairwise comparisons as a graph. Inspired by abend-etal-2015-lexical, who used combinatorial optimisation techniques to recover the lexical order of recipe pieces (abend-etal-2015-lexical), we propose Travelling Salesman Problem Rank (TSPRank). This hybrid pairwise-listwise method reframes the ranking problem as a Travelling Salesman Problem (TSP), an NP-hard combinatorial optimisation challenge that seeks the optimal sequence of node visits to minimise total travel cost. In a ranking context, TSPRank aims to find the optimal permutation of ranking entities to maximise the ranking score.
@@ 2. Background
LETOR algorithms aim to rank entities based on their features. In LETOR, we denote a list of $N$ ranking entities as $\mathbf{e}$, containing $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$, with each $\mathbf{e}_{i}\in\mathbb{R}^{d}$ where $d$ represents the dimensionality of the feature space. A scoring function $s:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is applied on the entities, which are then ranked in descending order of the $s(\mathbf{e}_{i})$ scores. LETOR algorithms are broadly categorised into pointwise, pairwise, and listwise approaches based on their optimisation strategies. Pointwise methods treat ranking as a regression problem, comparing scores $s(\mathbf{e}_{i})$ to labels $y_{i}$. Models such as OPRF (fuhr1989optimum), TreeBoost (friedman2001greedy), and RankSVM (shashua2002taxonomy), including newer implementations like RankCNN (severyn2015learning), are known for their computational efficiency but may not capture inter-entity relationships adequately.
@@
Pairwise methods focus on the relative comparisons of entity pairs, assessing scores $(s(\mathbf{e}_{i}),s(\mathbf{e}_{j}))$ against binary labels $y_{ij}$, which denote if entity $\mathbf{e}_{i}$ should be ranked higher than $\mathbf{e}_{j}$. Methods in this category include RankNet (burges2005learning), LambdaRank (burges2006learning), and LambdaMART (lambdamart). Listwise methods directly compute scores based on the features of all entities in a list. These methods aim to optimise a listwise objective, comparing the entire set of computed scores $s(\mathbf{e})\in\mathbb{R}^{N}$ against a complete set of labels $\mathbf{y}\in\mathbb{R}^{N}$. To advance this approach, aiLearningDeepListwise2018a introduced the Deep Listwise Context Model (DLCM) (aiLearningDeepListwise2018a). This model was further refined into general multivariate scoring functions in (aiLearningGroupwiseMultivariate2019). Considering that a list of entities can be viewed as a sequence, methodologies from Natural Language Processing (NLP) are applicable to ranking tasks. Notably, the Transformer architecture (attention), has been adapted to address listwise ranking challenges, referred to as listwise Transformers in recent work (peiPersonalizedRerankingRecommendation2019; pasumarthiPermutationEquivariantDocument2020; pang2020setrank; zhuangCrossPositionalAttentionDebiasing2021; liPEARPersonalizedReranking2022; rankformer).
@@
Although listwise ranking methods, primarily neural rankers, have demonstrated their advantages in web search reranking tasks (aiLearningDeepListwise2018a; pang2020setrank), google-neural-outperform-gbdt conducted a comprehensive evaluation comparing listwise neural rankers to the gradient-boosted-decision-tree (GBDT)-based pairwise LambdaMART (lambdamart; google-neural-outperform-gbdt). The results indicated that LambdaMART consistently outperforms listwise rankers across three widely used information retrieval datasets. Furthermore, it was observed that listwise rankers require substantial efforts to achieve only marginal improvements over LambdaMART. Consequently, it is evident that listwise ranking models are not universally applicable solutions and generally lack the robustness needed for broader deployment. This conclusion has also been further validated by rankformer (rankformer).
@@ 3. TSPRank
Directly predicting the listwise order is challenging, as correctly ranking $N$ entities from 1 to $N$ is complex. However, breaking it down into $N\times N$ pairwise comparisons simplifies the task, as each pairwise comparison is more straightforward than the entire listwise ranking. However, a fundamental drawback of pairwise ranking, as highlighted by pairwise-to-listwise (pairwise-to-listwise), is that the learning objectives do not focus on minimising the listwise ranking errors. To address this gap and rethink the approach of pairwise ranking, we introduce TSPRank, a novel methodology for position-based ranking problems. This approach integrates listwise optimisation into pairwise comparisons by modelling ranking tasks as TSP. Notably, TSPRank can be applied as an additional component with embeddings generated from domain-specific backbone models to achieve improved ranking performance.
@@ 3.2. Scoring Model
The scoring model takes a set of ranking entities or their corresponding embeddings $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$ as input, where each $\mathbf{e}_{i}\in\mathbb{R}^{d}$ and $d$ is the dimension of the features or embeddings. It outputs an $N\times N$ weighted adjacency matrix $A$ that represents the pairwise scores of every two entities. The scoring model comprises an optional node encoder and a trainable bilinear model on top.
@@
The default encoder used is the Transformer encoder block (attention), but it can be replaced with any other encoder suited to the specific task and dataset. Additionally, using an encoder is optional if sufficiently robust backbone models are available.
@@ 3.3. Inference
Inference of TSPRank, described in Equation 1, requires solving the open-loop TSP, which is widely recognised as an NP-hard combinatorial optimisation problem. The solution algorithms of the TSP consist of exact algorithms (brunch and bound (poikonen2019branch)) and approximation algorithms (ant colony optimisation (yang2008ant)). To ensure the desired ranking accuracy, we prefer to obtain an exact solution for the TSP rather than an approximate one. Therefore, we formulate the TSP as a Mixed Integer Linear Programming (MILP) problem (rieck2010new), where the integer variables represent binary decisions on whether an edge between two entities is included in the optimal tour. We then use an MILP optimisation solver, such as CPLEX (cplex) or Gurobi (gurobi), to find the exact optimal solution.
@@
(4) $x_{ij}=\begin{cases}1,&\text{if entity $\mathbf{e}_{j}$ is ranked immediately after $\mathbf{e}_{i}$,}\\ 0,&\text{otherwise.}\end{cases}$
@@ 4. Learning
We introduce two distinct learning methodologies for the TSPRank model outlined in Section 3. The first method is a local approach, focusing exclusively on individual pairwise comparisons during the training phase. The second method is a global, end-to-end approach, incorporating the Gurobi solver directly into the training process, thereby allowing the black-box MILP solver to influence the learning dynamically (li-etal-2023-bert). Figure 1 illustrates the complete pipeline of the two learning methods as well as the model architecture.
@@ 4.1. Local Learning
In the local learning approach, the objective is to understand pairwise consecutive relationships between entities. The objective is to determine whether entity $\mathbf{e}_{j}$ should rank one position after $\mathbf{e}_{i}$ in a given pair of nodes. Therefore, the score $s(\mathbf{e}_{i},\mathbf{e}_{j})$ is set high if $\mathbf{e}_{j}$ is supposed to rank consecutively after $\mathbf{e}_{i}$, and low otherwise.
@@ 4.2. Global Learning
(7) $\mathcal{L}_{glob}(A^{p},\mathbf{x}^{t})$ $=\max(0,\max_{\mathbf{x}\neq\mathbf{x}^{t}}[\Delta(\mathbf{x},\mathbf{x}^{t})+\mathbf{x}\cdot A^{p}]-\mathbf{x}^{t}\cdot A^{p}).$
The structured score $\Delta(\mathbf{x},\mathbf{x}^{t})$, defined in Equation 8, aims to enforce a margin for each incorrectly identified edge:
@@ 5.1. Benchmark Models
We benchmark our pairwise-listwise TSPRank model against the original backbone models (if they originally contain a prediction layer), LambdaMART (lambdamart), and Rankformer (rankformer), as LambdaMART is still considered the SOTA pairwise ranking method and Rankformer is the SOTA deep learning-based listwise ranking method. The benchmark models chosen include the best pure pairwise and pure listwise methods, which is sufficient to demonstrate the effectiveness of our approach.
@@ 5.2. Datasets
The task of stock ranking focuses on accurately predicting and ranking stocks according to their anticipated future returns ratio, which aids investors in selecting stocks for investment purposes (temporal-stock-prediction; zhang2022constructing). We use the dataset [^1] https://github.com/fulifeng/Temporal_Relational_Stock_Ranking/ introduced by temporal-stock-prediction (temporal-stock-prediction), which includes historical trading data from 2013 to 2017 for two significant markets, NASDAQ and NYSE, containing 1026 and 1737 stocks, respectively. We maintain a consistent setting for splitting the data into training, validation, and testing sets over a 3-year, 1-year, and 1-year period.
@@
Information retrieval is another critical area where ranking models are extensively applied. The task of information retrieval focuses on accurately ranking documents based on their relevance to a given query, which is crucial for search engines. LambdaMART and Rankformer are both originally proposed for information retrieval task (lambdamart; rankformer). We use the MQ2008-list [^2] https://www.microsoft.com/en-us/research/project/letor-learning-rank-information-retrieval/letor-4-0/ dataset from Microsoft LETOR4.0 (letor4), a popular benchmark dataset for LETOR algorithms. We acknowledge the existence of more recent information retrieval datasets, but they only provide relevance-level labels without ordinal ranking labels, making them unsuitable for our model evaluation.
@@
In the final task, we include a textual dataset called “On This Day 2” (OTD2) [^3] https://github.com/ltorroba/machine-reading-historical-events for chronologically ordering historical events. Originally, honovich-etal-2020-machine constructed this dataset for a regression task to predict the year of occurrence (honovich-etal-2020-machine). However, it can also be approached as a ranking task to predict the chronological order of events within a given group of events. The OTD2 dataset, sourced from the “On This Day” website, includes 71,484 events enriched with additional contextual information. Scraped in April 2020, it incorporates recent events and corrections by the site. Preprocessing excluded events before 1 CE and future projections such as “31st predicted perihelion passage of Halley’s Comet” in 2061 CE. We follow the 80%/10%/10% train/validation/test split as provided.

### v3

@@ Abstract.
Traditional Learning-To-Rank (LETOR) approaches, including pairwise methods like RankNet and LambdaMART, often fall short by solely focusing on pairwise comparisons, leading to sub-optimal global rankings. Conversely, deep learning based listwise methods, while aiming to optimise entire lists, require complex tuning and yield only marginal improvements over robust pairwise models. To overcome these limitations, we introduce Travelling Salesman Problem Rank (TSPRank), a hybrid pairwise-listwise ranking method. TSPRank reframes the ranking problem as a Travelling Salesman Problem (TSP), a well-known combinatorial optimisation challenge that has been extensively studied for its numerous solution algorithms and applications. This approach enables the modelling of pairwise relationships and leverages combinatorial optimisation to determine the listwise ranking. This approach can be directly integrated as an additional component into embeddings generated by existing backbone models to enhance ranking performance. Our extensive experiments across three backbone models on diverse tasks, including stock ranking, information retrieval, and historical events ordering, demonstrate that TSPRank significantly outperforms both pure pairwise and listwise methods. Our qualitative analysis reveals that TSPRank’s main advantage over existing methods is its ability to harness global information better while ranking. TSPRank’s robustness and superior performance across different domains highlight its potential as a versatile and effective LETOR solution. The code and preprocessed data are available at https://github.com/waylonli/TSPRank-KDD2025.
@@ 1. Introduction
Learning to Rank (LETOR) algorithms have become essential in applications such as recommendation systems (Fan et al., 2022; Pei et al., 2019; Li et al., 2021), question answering (Ji and Wang, 2013; Campese et al., 2024), and information retrieval (Ai et al., 2018; Pang et al., 2020; Buyl et al., 2023). These algorithms aim to order a list of ranking entities based on their features, optimising for the most relevant or preferred entities to appear at the top. Recently, LETOR methods have also expanded into other domains, such as stock and portfolio selection (Song et al., 2017; Wu et al., 2024) and textual ordering (Tan et al., 2013; Wang et al., 2023). Despite these broader applications of LETOR, latest fundamental work on LETOR still primarily focused on incorporating click data into ranking models and addressing biases introduced by user feedback (Oosterhuis, 2023; Zhang et al., 2023; Guo et al., 2020), concentrating on retrieval and recommendation tasks. Limited research has explored general ranking methods across diverse tasks and domains.
@@
Over the past decade, research on LETOR models mainly focuses on pairwise and listwise approaches, while pointwise methods often fail to capture the intricate inter-entity relationships that are essential for accurate ranking. Pairwise methods, such as LambdaMART and RankNet (Burges, 2010), primarily optimise for pairwise comparisons without a holistic view of the entire ranking list, potentially leading to sub-optimal global rankings. Listwise methods optimise the ranking of entire lists directly rather than individuals or pairs. This category includes advanced neural network architectures, particularly adaptations of the Transformer architecture (Vaswani et al., 2017), such as Rankformer (Buyl et al., 2023) and SetRank (Pang et al., 2020). However, existing work shows that deep learning-based listwise models require complex tuning to achieve marginal gains over robust pairwise models like LambdaMART on information retrieval benchmarks (Qin et al., 2021). Therefore, pairwise and listwise methods each have inherent drawbacks and more effective solutions for LETOR have not been exhaustively explored.
@@
Recently, GNNRank successfully used directed graph neural networks to learn listwise rankings from pairwise comparisons (He et al., 2022). Although their method is not applicable to general LETOR problems, as it requires pre-known pairwise relationships like the outcomes of sports matches, it still suggests a potential approach for LETOR by representing pairwise comparisons as a graph. Inspired by Abend et al., who used combinatorial optimisation techniques to recover the lexical order of recipe pieces (Abend et al., 2015), we propose Travelling Salesman Problem Rank (TSPRank). This hybrid pairwise-listwise method reframes the ranking problem as a Travelling Salesman Problem (TSP), an NP-hard combinatorial optimisation challenge that seeks the optimal sequence of node visits to minimise total travel cost. In a ranking context, TSPRank aims to find the optimal permutation of ranking entities to maximise the ranking score.
@@ 2. Background
LETOR algorithms aim to rank entities based on their features. In LETOR, we denote a list of $N$ ranking entities as $\mathbf{e}$, containing $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$, with each $\mathbf{e}_{i}\in\mathbb{R}^{d}$ where $d$ represents the dimensionality of the feature space. A scoring function $s:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is applied on the entities, which are then ranked in descending order of the $s(\mathbf{e}_{i})$ scores. LETOR algorithms are broadly categorised into pointwise, pairwise, and listwise approaches based on their optimisation strategies. Pointwise methods treat ranking as a regression problem, comparing scores $s(\mathbf{e}_{i})$ to labels $y_{i}$. Models such as OPRF (Fuhr, 1989), TreeBoost (Friedman, 2001), and RankSVM (Shashua and Levin, 2002), including newer implementations like RankCNN (Severyn and Moschitti, 2015), are known for their computational efficiency but may not capture inter-entity relationships adequately.
@@
Pairwise methods focus on the relative comparisons of entity pairs, assessing scores $(s(\mathbf{e}_{i}),s(\mathbf{e}_{j}))$ against binary labels $y_{ij}$, which denote if entity $\mathbf{e}_{i}$ should be ranked higher than $\mathbf{e}_{j}$. Methods in this category include RankNet (Burges et al., 2005), LambdaRank (Burges et al., 2006), and LambdaMART (Burges, 2010). Listwise methods directly compute scores based on the features of all entities in a list. These methods aim to optimise a listwise objective, comparing the entire set of computed scores $s(\mathbf{e})\in\mathbb{R}^{N}$ against a complete set of labels $\mathbf{y}\in\mathbb{R}^{N}$. To advance this approach, Ai et al. introduced the Deep Listwise Context Model (DLCM) (Ai et al., 2018). This model was further refined into general multivariate scoring functions in (Ai et al., 2019). Considering that a list of entities can be viewed as a sequence, methodologies from Natural Language Processing (NLP) are applicable to ranking tasks. Notably, the Transformer architecture (Vaswani et al., 2017), has been adapted to address listwise ranking challenges, referred to as listwise Transformers in recent work (Pei et al., 2019; Pasumarthi et al., 2020; Pang et al., 2020; Zhuang et al., 2021; Li et al., 2022; Buyl et al., 2023).
@@
Although listwise ranking methods, primarily neural rankers, have demonstrated their advantages in web search reranking tasks (Ai et al., 2018; Pang et al., 2020), Qin et al. conducted a comprehensive evaluation comparing listwise neural rankers to the gradient-boosted-decision-tree (GBDT)-based pairwise LambdaMART (Burges, 2010; Qin et al., 2021). The results indicated that LambdaMART consistently outperforms listwise rankers across three widely used information retrieval datasets. Furthermore, it was observed that listwise rankers require substantial efforts to achieve only marginal improvements over LambdaMART. Consequently, it is evident that listwise ranking models are not universally applicable solutions and generally lack the robustness needed for broader deployment. This conclusion has also been further validated by Buyl et al. (Buyl et al., 2023).
@@ 3. TSPRank
Directly predicting the listwise order is challenging, as correctly ranking $N$ entities from 1 to $N$ is complex. However, breaking it down into $N\times N$ pairwise comparisons simplifies the task, as each pairwise comparison is more straightforward than the entire listwise ranking. However, a fundamental drawback of pairwise ranking, as highlighted by Cao et al. (Cao et al., 2007), is that the learning objectives do not focus on minimising the listwise ranking errors. To address this gap and rethink the approach of pairwise ranking, we introduce TSPRank, a novel methodology for position-based ranking problems. This approach integrates listwise optimisation into pairwise comparisons by modelling ranking tasks as TSP. Notably, TSPRank can be applied as an additional component with embeddings generated from domain-specific backbone models to achieve improved ranking performance.
@@ 3.2. Scoring Model
The scoring model takes a set of ranking entities or their corresponding embeddings $\{\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{N}\}$, where each $\mathbf{e}_{i}\in\mathbb{R}^{d}$ and $d$ is the dimension of the features or embeddings. It outputs an $N\times N$ weighted adjacency matrix $A$ that represents the pairwise scores of every two entities. The scoring model comprises an optional node encoder and a trainable bilinear model on top.
@@
The default encoder used is the Transformer encoder block (Vaswani et al., 2017), but it can be replaced with any other encoder suited to the specific task and dataset. Additionally, using an encoder is optional if sufficiently robust backbone models are available.
@@ 3.3. Inference
Inference of TSPRank, described in Equation 1, requires solving the open-loop TSP, which is widely recognised as an NP-hard combinatorial optimisation problem. The solution algorithms of the TSP consist of exact algorithms (brunch and bound (Poikonen et al., 2019)) and approximation algorithms (ant colony optimisation (Yang et al., 2008)). Given the scale of the ranking problem and the desired ranking accuracy, we prefer to obtain an exact solution for the TSP rather than an approximate one. Therefore, we formulate the TSP as a Mixed Integer Linear Programming (MILP) problem (Rieck and Zimmermann, 2010), where the integer variables represent binary decisions on whether a path between two nodes is included in the optimal tour. We then use an MILP optimization solver, such as CPLEX (Cplex, 2009) or Gurobi (Gurobi Optimization, LLC, 2023), to find the exact optimal solution.
@@
(4) $x_{ij}=\begin{cases}1,&\text{if entity $\mathbf{e}_{j}$ is ranked immediately after $\mathbf{e}_{i}$.}\\ 0,&\text{otherwise.}\end{cases}$
@@ 4. Learning
We introduce two distinct learning methodologies for the TSPRank model outlined in Section 3. The first method is a local approach, focusing exclusively on individual pairwise comparisons during the training phase. The second method is a global, end-to-end approach, incorporating the Gurobi solver directly into the training process, thereby allowing the black-box MILP solver to influence the learning dynamically (Li et al., 2023). Figure 1 illustrates the complete pipeline of the two learning methods as well as the model architecture.
@@ 4.1. Local Learning
In the local learning approach, the objective is to understand pairwise consecutive relationships between nodes. The objective is to determine whether entity $\mathbf{e}_{j}$ should rank one position after $\mathbf{e}_{i}$ in a given pair of nodes. Therefore, the score $s(\mathbf{e}_{i},\mathbf{e}_{j})$ is set high if $\mathbf{e}_{j}$ is supposed to rank consecutively after $\mathbf{e}_{i}$, and low otherwise.
@@ 4.2. Global Learning
(7) $\mathcal{L}_{glob}(A^{p},\mathbf{x}^{t})$ $=\max(0,\max_{\mathbf{x}\neq\mathbf{x}^{t}}[\Delta(\mathbf{x},\mathbf{x}^{t})+\mathbf{x}\cdot A^{p}]-\mathbf{x}^{t}\cdot A^{p})$
The structured score $\Delta(\mathbf{x},\mathbf{x}^{t})$, defined in Equation 8, aims to enforce a margin for each incorrectly identified edge.
@@ 5.1. Benchmark Models
We benchmark our pairwise-listwise TSPRank model against the original backbone models (if they originally contain a prediction layer), LambdaMART (Burges, 2010), and Rankformer (Buyl et al., 2023), as LambdaMART is still considered the SOTA pairwise ranking method and Rankformer is the SOTA deep learning-based listwise ranking method. The benchmark models chosen include the best pure pairwise and pure listwise methods, which is sufficient to demonstrate the effectiveness of our approach.
@@ 5.2. Datasets
The task of stock ranking focuses on accurately predicting and ranking stocks according to their anticipated future returns ratio, which aids investors in selecting stocks for investment purposes (Feng et al., 2019; Zhang et al., 2022). We use the dataset [^1] https://github.com/fulifeng/Temporal_Relational_Stock_Ranking/ introduced by Feng et al. (Feng et al., 2019), which includes historical trading data from 2013 to 2017 for two significant markets, NASDAQ and NYSE, containing 1026 and 1737 stocks, respectively. We maintain a consistent setting for splitting the data into training, validation, and testing sets over a 3-year, 1-year, and 1-year period.
@@
Information retrieval is another critical area where ranking models are extensively applied. The task of information retrieval focuses on accurately ranking documents based on their relevance to a given query, which is crucial for search engines. LambdaMART and Rankformer are both originally proposed for information retrieval task (Burges, 2010; Buyl et al., 2023). We use the MQ2008-list [^2] https://www.microsoft.com/en-us/research/project/letor-learning-rank-information-retrieval/letor-4-0/ dataset from Microsoft LETOR4.0 (Qin and Liu, 2013), a popular benchmark dataset for LETOR algorithms. We acknowledge the existence of more recent information retrieval datasets, but they only provide relevance-level labels without ordinal ranking labels, making them unsuitable for our model evaluation.
@@
In the final task, we include a textual dataset called “On This Day 2” (OTD2) [^3] https://github.com/ltorroba/machine-reading-historical-events for chronologically ordering historical events. Originally, Honovich et al. constructed this dataset for a regression task to predict the year of occurrence (Honovich et al., 2020). However, it can also be approached as a ranking task to predict the chronological order of events within a given group of events. The OTD2 dataset, sourced from the “On This Day” website, includes 71,484 events enriched with additional contextual information. Scraped in April 2020, it incorporates recent events and corrections by the site. Preprocessing excluded events before 1 CE and future projections such as “31st predicted perihelion passage of Halley’s Comet” in 2061 CE. We follow the 80%/10%/10% train/validation/test split as provided.